OpenStudy (anonymous):

Jon is building a dog run up against his house. He has 64 feet of fencing to use on the other three sides. What would be the maximum area possible for his dog run? a)96ftsqd b)144ftsqd c)256ftsqd d)512ftsqd

OpenStudy (amistre64):

a square is the largest area rectangle so; (64/3)^2 = 455 but that aint one of your options is it....hmmm

OpenStudy (amistre64):

3x = 64 x = 64/3 Area = x*x (64/3)*(64/3) = 455. hmmm

OpenStudy (amistre64):

you sure you aint got a typo?

OpenStudy (anonymous):

i dont think so :(

OpenStudy (amistre64):

the biggest area you can get from 4 sided figures with equal perimeters is a square... so i dont really know what they are asking; unless they want to kno wwhich one of these will fit

OpenStudy (amistre64):

but that doesnt fit the question asked does it; the "maximum area possible" doesnt mean "which of these is the max area you could make from it"

OpenStudy (anonymous):

i guess i mightve typoed but i dont have the sheet to look @ with me, thnks tho

OpenStudy (amistre64):

youre welcome; wish i could have been more helpful on it still :)